How do I solve this equation?
9+3/x-3=x/x-3

Answer by stanbon(57246)

solve this equation?
9 + (3/x-3) = (x/x-3)
lcm: x-3
Multiply thru by the lcm to get:
9(x-3) + 3 = x
9x - 27 + 3 = x
8x = 24
x = 3
====================
But x cannot be 3 in the original equation as that would make the denominators zero.
Conclusion: No solution
------------------------
Cheers,
Stan H.